# Online Multivariate Regularized Distributional Regression for High-dimensional Probabilistic Electricity Price Forecasting

### v1

@@ 2.4 Online Estimation Algorithm
High-level Overview
The IRLS algorithm consists of two nested loops. In the outer loop, we iterate through all distribution parameters. In the inner loop, we repeatedly run a weighted fit of the score vector $\bm{\mathrm{u}}$ on the design matrix $\bm{\mathrm{X}}$ using the weights $\bm{\mathrm{W}}$ until convergence. Note that in the inner loop, we run the weighted fit sequentially for all elements of the distribution parameter. Since the fit itself is agnostic to the regression technique (Stasinopoulos et al., 2024), we employ the online coordinate descent-based LASSO estimation here, as it has been proposed by Hirsch et al. (2024) for the univariate case already. Algorithm 2 gives an overview on the online estimation of multivariate distributional regression models. We define the index sets $\mathcal{K}=\{0,1,...,p-1\}$ for the number of parameters and $\mathcal{M}_{k}=\{0,1,..,M_{k}-1\}$ for the number of elements of each parameter as described in Section 2.1.
@@
Update of the Gramian and Weights
For each inner iteration $i$, the update of the Gramian matrices starts at the Gramian matrices of $\bm{\mathrm{G}}_{km}^{[n]}$ and $\bm{\mathrm{H}}_{km}^{[n]}$ and the new information enters the Gramian matrices through the update of the weights and the working vector. However, the weights are also updated iteratively along each inner and outer iteration $i$ and $r$ due to the Newton-Raphson step towards the optimal coefficients. The weights can only be updated for the current update step $n+1$, while previous weights remain fixed. In a pure batch case, all weights are updated within each Newton-Raphson step. This introduces an approximation error for the online case, which can be controlled by the forget parameter $\gamma$ as shown in Hirsch et al. (2024).
Model selection
@@
where $\ell$ is the log-likelihood under the model, $K$ is the number of parameters in the model and $N$ the number of seen observations. Let us note that we can recover Akaikes Information Criterion (AIC), the Bayesian Information Criterion (BIC) and the Hannan-Quinn Information Criterion (HQC) by setting $\nu_{0},\nu_{1},\nu_{2}$ accordingly. The optimal regularization parameter is then selected as $\lambda_{mk}^{\text{opt}}=\operatorname{argmin}_{\lambda}\operatorname{IC}$. Since the evaluation of the likelihood can be costly for high-dimensional data, we propose to employ the first derivative of the log-likelihood, i.e. calculate
@@
where the superscript ${[\lambda_{i}]}$ denotes the model with the regularization parameter $\lambda_{i}$. The approximation is valid for small changes in the regularization parameter and avoids the costly re-evaluation of the likelihood.
Step-size and Damping
The algorithm goes iteratively along all coordinates of the distribution parameter. The coordinates of the distribution parameters might impact each other, e.g. in the matrix multiplication of the CD-based scale matrix (see also the definition of the derivatives in A.4 and A.5). At the same time, we initialize the fitted values $\hat{\theta}_{m}$ as constant values. To stabilize the estimation, we propose to update the values in the very first iteration $i$ by a “dampened” version, i.e. taking
@@
Parallelization
Since the partial derivatives are not information orthogonal, the options for parallelization remain limited unfortunately. For the multivariate normal and $t$-distribution used in this paper, only the estimation of the location parameter can be parallelized in any case, as well as the estimation of the coordinates of the LRA matrix $\bm{\mathrm{A}}=\operatorname{diag}(a_{1},...,a_{D})$ for the normal distribution. For the $t$-distribution, the estimation of $\bm{\mathrm{A}}$ can only be parallelized for sufficiently high degrees of freedom. As parallelization would incur further open questions with respect to individual or joint regularization and model selection and the location parameter generally converges rather fast, we have not implemented parallel computation yet.
@@ 2.5 Path-based Regularized Estimation for the Scale Matrix
Idea
Often, some structure can be imposed on the covariance matrix, i.e. in spatial or temporal data, which has a clear dependence pattern along the diagonal. In these cases, the covariance matrix can be regularized by systematically setting far off-diagonal elements to zero (Gabriel, 1962; Zimmerman and Núñez-Antón, 1997; Zimmerman et al., 1998). While both, the CD-based and the LRA-based scale matrix parametrization lend themselves to this type of regularization, the approach is mainly popular with the Cholesky-based parameterization due to the relationship between the elements of the CD and the temporal correlation for longitudinal data under the name AD-$r$ regularization. However, such regularization is commonly applied a-priori and not in a data-driven fashion, see e.g. Muschinski et al. (2022); Zimmerman and Núñez-Antón (1997). On the other side, in coordinate descent estimation of regularized problems such as LASSO, path-based estimation starting from a strongly regularized solution towards an (almost) not regularized solution has proven itself as efficient solution approach. In this section, we aim to combine these two principles by introducing path-based estimation for the regularized scale matrix. On a high level, our algorithm starts with an “independence-parameterization” of the scale matrix and subsequently adds more non-zero elements and thus complexity to the parameterization of the scale matrix. Figure 4 illustrates how the path-based estimation uses increasingly complex specifications for the scale matrix $\bm{\mathrm{\Sigma}}$ respectively $\bm{\mathrm{\Omega}}$. Formally, for some regularization parameter $\alpha$, we set the elements of the scale to zero for
  • the CD-based parameterization if the indices $i,j$ are such that $|i-j|>\alpha$,
  • the LRA-based parameterization if the indices $d,r$ of $\bm{\mathrm{V}}$ are such that $r\geq\alpha$
@@
Figure 4: Path-based estimation along increasingly complex scale matrix parameterizations. The top panel shows the AD-$r$ regression for a Cholesky-based parameterization. The lower panel shows the estimation along the LRA-based parameterization, where $\bm{\mathrm{A}}=\operatorname{diag}(a_{1},...,a_{d})$ is not regularized and the $D\times r$ matrix $\bm{\mathrm{V}}$ is filled column-wise with non-zero elements. Own Illustration.
1 for $\alpha=0,...,D$ do
@@
  • For a (small) fixed maximum regularization size, the number of parameters in the CD-based parameterization grows (almost) linear in $D$, alleviating the disadvantage of quadratic complexity.
  • For the multivariate $t$-distribution, independence is only achieved as $\nu\rightarrow\infty$. We therefore set a high initial guess ($\nu=10\operatorname{e}^{6}$) for the first outer iteration of $\bm{\mathrm{\mu}}$ and $\bm{\mathrm{\Omega}}$ to ensure numerical stability for the first iteration and subsequently choose a lower initial guess for the first iteration of $\nu$, since the Newton-Raphson algorithm relies on appropriate start values and tends to alternate between extrema otherwise (see e.g. Casella and Bachmann, 2021; Kornerup and Muller, 2006, on the impact of initial values for Newton-Raphson algorithms.). [^1] We have found the algorithm to iterate between $\nu=2$ and $\nu>10\text{e}^{10}$ for too large start values for the degrees of freedom. The proposed approach however has proved stable through the full simulation study with highly volatile electricity prices.
  • We can employ the path-based estimation to early stop the estimation, if the log-likelihood respectively an information criterion does not increase sufficiently by adding more non-zero elements. This allows for both, implicit regularization and decreased estimation time. However, once we early stop, we cannot increase the complexity of the parametrization in the online estimation but need to treat this as fixed.
Currently, the Algorithm will add only full off-diagonals (CD) respectively columns (LRA). The implementation however could also work for block-wise schemes (see e.g. the adaptive block structure in Cai and Yuan, 2012) or user-defined regularization patterns. The development of smart selection schemes for the next coordinates of the covariance matrix to include would be much beneficial for the speed of the algorithm.
@@ 3 Forecasting Study
Electricity Market Design
For electricity produced on day $t$ and hour $h$, the short-term electricity market in Germany is split in three major parts: The daily day-ahead auction on $t-1$ at 12:00 hours for 24 hourly delivery periods $h\in\{0,...,23\}$, the afternoon auction with quarter-hourly delivery periods on $t-1$, at 15:00 hours and the continuous intraday market. In 2024, two additional 15-minute auctions have been introduced at $t-1$, 22:00 hours and 10:00 on $t$. The daily procedure for the day-ahead auction, which is the focus of this paper, is shown in Figure 5. The market is organized by EPEX SPOT and Nordpool in the joint single day-ahead coupling (SDAC) as a pay-as-cleared auction through the EUPHEMIA algorithm, resembling the merit-order model for the electricity market (Billé et al., 2023; Hirsch et al., 2024; Viehmann, 2017).
[…]
@@
[…]
@@
Online Forecasting Study
At each day $t$, before the day-ahead auction at 12:00, we aim to generate forecasts for day $t+1$. Prior to forecasting, we update our models by taking into account the realized prices and forecasts for delivery day $t$. Figure 6 shows the structure of the online forecasting study. We are among the first studies to enforce a strict online setting for the forecasting study. Hence, on each day $t$, models are only updated using information revealed to the forecaster on this day. This is in contrast to the repeated batch learning approach, where the model is re-estimated on the full data set after each day.
Data and Notation
We use the same data set as in Marcjasz et al. (2023); Hirsch et al. (2024); Brusaferri et al. (2024a), which consists of electricity prices for the German day-ahead market from 2015-01-01 to 2021-01-01. In line with previous works, we use the data until 2018-12-26 as initial training set, leaving 736 observations and therefore more than 2 years, as it is best practice (Lago et al., 2021), for out-of-sample testing. Additionally, the data set contains day-ahead renewable production forecasts, load forecasts and prices for fundamental commodities. The data is available at the Github repository of Marcjasz et al. (2023). We denote the electricity price for day $t$ and hour $h\in 0,...,23$ as $y_{th}$ and therefore have $\bm{\mathrm{Y}}=(\bm{\mathrm{y}}_{0},...,\bm{\mathrm{y}}_{23})$ as the 24-dimensional ($D=H=24$) response matrix. We therefore, in this application study, we have $T$ corresponding to $N$ and $H$ corresponding to $D$ in the general notation. All fundamental features are briefly described in Table 2.
@@
$\operatorname{RES}_{t,h}$ Day-ahead Renewable Energy Production Forecast Hourly ENTSO-E
$\overline{\operatorname{RES}}_{t}$ Day-ahead baseload RES Forecast $\frac{1}{H}\sum_{h=1}^{H}\operatorname{RES}_{t,h}$ Daily ENTSO-E
$\operatorname{Load}_{t,h}$ Day-ahead Electricity Load Forecast Hourly ENTSO-E
$\overline{\operatorname{Load}}_{t}$ Day-ahead baseload Load Forecast $\frac{1}{H}\sum_{h=1}^{H}\operatorname{Load}_{t,h}$ Daily ENTSO-E
@@
Table 2: Variables from the data set of Marcjasz et al. (2023).
Model definition

### v2

@@ 2.5 Online Estimation Algorithm
As outlined in Section 2.3 and Algorithm 1, the IRLS algorithm consists of two nested loops. In the outer loop, we iterate through all distribution parameters. In the inner loop, we repeatedly run a weighted fit of the score vector $\bm{\mathrm{u}}$ on the design matrix $\bm{\mathrm{X}}$ using the weights $\bm{\mathrm{W}}$ until convergence. Note that in the inner loop, we run the weighted fit sequentially for all elements of the distribution parameter. Since the fit itself is agnostic to the regression technique (Stasinopoulos et al., 2024), we employ the online coordinate descent-based LASSO estimation here, as it has been proposed by Hirsch et al. (2024) for the univariate case already. Algorithm 3 gives an overview on the online estimation of multivariate distributional regression models. While its general structure follows the high-level overview in Algorithm 1, we provide a detailed treatment on the online update using OCD and the implications due to the multivariate case. We define the index sets $\mathcal{K}=\{1,...,p\}$ for the number of parameters and $\mathcal{M}_{k}=\{1,..,M_{k}\}$ for the number of elements of each parameter as described in Section 2.4.
@@
For each inner iteration $i$, the update of the Gramian matrices starts at the Gramian matrices of $\bm{\mathrm{G}}_{km}^{[n]}$ and $\bm{\mathrm{h}}_{km}^{[n]}$ and the new information enters the Gramian matrices through the update of the weights $\bm{\mathrm{W}}$ and the working vector $\bm{\mathrm{z}}$. However, the weights are also updated iteratively along each inner and outer iteration $i$ and $r$ due to the Newton-Raphson step towards the optimal coefficients. The weights can only be updated for the current update step $n+1$, while previous weights remain fixed. In a pure batch case, all weights are updated within each Newton-Raphson step. This introduces an approximation error for the online case, which can be controlled by the forget parameter $\gamma$ as shown in Hirsch et al. (2024).
@@
where $\ell$ is the log-likelihood under the model, $K$ is the number of parameters in the model and $N$ the number of seen observations. We can recover Akaikes Information Criterion (AIC), the Bayesian Information Criterion (BIC), and the Hannan-Quinn Information Criterion (HQC) by setting $\nu_{0},\nu_{1},\nu_{2}$ accordingly. The optimal regularization parameter is then selected as $\lambda_{mk}^{\text{opt}}=\operatorname{argmin}_{\lambda}\operatorname{IC}$. Since the evaluation of the likelihood can be costly for high-dimensional data, we propose to employ the first derivative of the log-likelihood, i.e. calculate
@@
where the superscript $\lambda_{i}]$} denotes the model with the regularization parameter $\lambda_{i}$. The approximation is valid for small changes in the regularization parameter and avoids the costly re-evaluation of the likelihood.
The algorithm goes iteratively along all coordinates of the distribution parameter. The coordinates of the distribution parameters might impact each other, e.g. in the matrix multiplication of the CD-based scale matrix (see also the definition of the derivatives in A.3 and A.4). At the same time, we initialize the fitted values $\hat{\theta}_{m}$ as constant values. To stabilize the estimation, we propose to update the values in the very first iteration $i$ by a “dampened” version, i.e. taking
@@
Since the partial derivatives are not information orthogonal, the options for parallelization remain limited unfortunately. For the multivariate normal and $t$-distribution used in this paper, only the estimation of the location parameter can be parallelized, as well as the estimation of the coordinates of the LRA matrix $\bm{\mathrm{A}}=\operatorname{diag}(a_{1},...,a_{D})$ for the normal distribution. For the $t$-distribution, the estimation of $\bm{\mathrm{A}}$ can only be parallelized for sufficiently high degrees of freedom. One plausible option for parallelization with non-orthogonal parameters would be using a step size smaller than one, i.e. using a convex combination of the previous and the newly estimated coefficients. However, this would introduce an additional hyperparameter. As parallelization would incur further open questions with respect to individual or joint regularization and model selection and the location parameter generally converges rather fast, we have not implemented parallel computation yet and leave it for future research.
@@ 2.6 Path-based Regularized Estimation for the Scale Matrix
Using LASSO-type regularization, the algorithm can handle high-dimensional covariate spaces for each coordinate of the distribution parameter. However, for high-dimensional response variables, i.e. large $D$, the number of parameters in the scale matrix grows quadratically in $D$ for the CD-based parameterizations and linearly in $D$ for the LRA-based parameterization (see Table 2). To alleviate this issue and allow for parsimonious modelling for large $D$, we propose a path-based estimation approach that starts with a simple (highly regularized) structure of the scale matrix and gradually increases its complexity. This approach is inspired by regularization techniques in time series analysis and path-based estimation methods in high-dimensional statistics (such as the graphical LASSO, Friedman et al., 2008). We exploit that in many cases, some structure can be imposed on the scale matrix, i.e. in spatial or temporal data, which has a clear dependence pattern along the diagonal. In these cases, the scale matrix can be regularized by systematically setting off-diagonal elements to zero (Gabriel, 1962; Zimmerman and Núñez-Antón, 1997; Zimmerman et al., 1998). While both, the CD-based and the LRA-based scale matrix parametrization lend themselves to this type of regularization, the approach is mainly popular with the Cholesky-based parameterizations due to the relationship between the elements of the CD and the temporal correlation for longitudinal data under the name AD-$r$ regularization. However, such regularization is commonly applied a-priori and not in a data-driven fashion, see e.g. Muschinski et al. (2022); Zimmerman and Núñez-Antón (1997). On the other side, in coordinate descent estimation of regularized problems such as (graphical) LASSO, path-based estimation starting from a strongly regularized solution towards an (almost) not regularized solution has proven itself as an efficient solution approach. In this section, we aim to combine these two principles by introducing path-based estimation for the regularized scale matrix.
On a high level, our algorithm starts with an “independence-parameterization” of the scale matrix and subsequently adds more non-zero elements and thus complexity to the parameterization of the scale matrix. Figure 4 illustrates how the path-based estimation uses increasingly complex specifications for the scale matrix $\bm{\mathrm{\Sigma}}$ respectively $\bm{\mathrm{\Omega}}$. Formally, for some regularization parameter $\alpha$, we set the elements of the scale to zero for
  • 1.
    the Cholesky-based parameterizations if the indices of the diagonal matrices $\bm{\mathrm{A}}$ or $\bm{\mathrm{L}}$, $i,j$ are such that $|i-j|>\alpha$,
  • 2.
    the LRA-based parameterization if the indices $d,r$ of $\bm{\mathrm{V}}$ are such that $r\geq\alpha$,
@@
Figure 4: Path-based estimation along increasingly complex scale matrix parameterizations. The top panel shows the AD-$r$ regression for a Cholesky-based parameterization $((\bm{\mathrm{L}}^{[\alpha]})^{-1})^{\top}(\bm{\mathrm{L}}^{[\alpha]})^{-1}$. The superscript $\alpha]$ denotes the iteration regularization. Black elements are the state after the initial fit assuming independence. Red elements of the vectors are added in iteration $\alpha$. Blue elements have been added in previous iterations. The lower panel shows the estimation along the LRA-based parameterization $\bm{\mathrm{A}}+(\bm{\mathrm{V}}^{[\alpha]})(\bm{\mathrm{V}}^{[\alpha]})^{\top}$, where $\bm{\mathrm{A}}=\operatorname{diag}(a_{1},...,a_{d})$ is not regularized and the $D\times r$ matrix $\bm{\mathrm{V}}$ is filled column-wise with non-zero elements. Own Illustration.
1for $\alpha=0,...,D$ do
@@
  • 1.
    For a (small) fixed maximum regularization size, the number of parameters in the CD and MCD-based parameterization grows (almost) linearly in $D$, alleviating the disadvantage of quadratic complexity.
  • 2.
    For the multivariate $t$-distribution, independence is only achieved as $\nu\rightarrow\infty$. We therefore set a high initial guess ($\nu=10^{6}$) for the first outer iteration of $\bm{\mathrm{\mu}}$ and $\bm{\mathrm{\Omega}}$ to ensure numerical stability for the first iteration and subsequently choose a lower initial guess for the first iteration of $\nu$, since the Newton-Raphson algorithm relies on appropriate start values and tends to alternate between extrema otherwise (see e.g. Casella and Bachmann, 2021; Kornerup and Muller, 2006, on the impact of initial values for Newton-Raphson algorithms.) [^2] We have found the algorithm to iterate between $\nu=2$ and $\nu>10^{10}$ for too large start values for the degrees of freedom. The proposed approach however has proved stable through the full simulation study with highly volatile electricity prices.
  • 3.
    We can employ the path-based estimation to early stop the estimation, if the log-likelihood or an information criterion does not increase sufficiently by adding more non-zero elements. This allows for both, implicit regularization and decreased estimation time. However, once we early stop, we cannot increase the complexity of the parametrization in the online estimation but need to treat this as fixed.
Currently, the Algorithm will add only full off-diagonals (for Cholesky-based approaches) respectively columns (LRA). The implementation however could also work for block-wise schemes (see e.g. the adaptive block structure in Cai and Yuan, 2012) or user-defined regularization patterns. The development of smart selection schemes for the next coordinates of the covariance matrix to include would be beneficial for the speed of the algorithm. We provide an analysis of the in-sample selection of $\alpha$ and the out-of-sample performance for various $\alpha$ in the forecasting study in Section 4.3 and leave the development of more advanced selection schemes for future research.
@@ 3 Forecasting Study
3.1 Day-ahead Electricity Market and Data
For electricity produced on day $t$ and hour $h$, the short-term electricity market in Germany is split in three major parts: The daily day-ahead auction on t1t-1 at 12:00 hours for 24 hourly delivery periods $h\in\{0,...,23\}$, the afternoon auction with quarter-hourly delivery periods on t1t-1, at 15:00 hours and the continuous intraday market. The daily procedure for the day-ahead auction, which is the focus of this paper, is shown in Figure 5. The market is organized by EPEX SPOT and Nordpool in the joint single day-ahead coupling (SDAC) as a pay-as-cleared auction through the EUPHEMIA algorithm, resembling the merit-order model for the electricity market (Billé et al., 2023; Hirsch et al., 2024; Viehmann, 2017).
Auction t1t-1, 12:00012Delivery periods h=0,,23h=0,...,23 on day tt222223Auction $t$, 12:00012Delivery periods h=0,,23h=0,...,23 on day t+1t+1222223Auction $t+1$, 12:00012Delivery periods h=0,,23h=0,...,23 on day t+2t+2222223Results at 12:42, t1t-1Results at 12:42, $t$
@@ 3.1 Day-ahead Electricity Market and Data
At each day $t$, before the day-ahead auction at 12:00, we aim to generate forecasts for day $t+1$. Prior to forecasting, we update our models by taking into account the realized prices and forecasts for delivery day $t$. Figure 6 contrasts online learning with the expanding and rolling window batch learning, two popular schemes for forecasting studies. Note that in the online learning scheme, we only use the new observation for updating the model, while in the batch learning schemes, we re-use all or a fixed window of previous observations. The rolling window scheme is popular in the EPF community (see e.g. Lago et al., 2021; Nowotarski and Weron, 2018) while online learning schemes are a rather recent development in EPF (e.g. Berrisch and Ziel, 2024; Zaffran et al., 2022; Hirsch et al., 2024).
Initial set i=0,,Ni=0,...,NBatch Expanding WindowDataset i=0,,N+1i=0,...,N+1Dataset i=0,,N+2i=0,...,N+2Dataset i=0,,N+3i=0,...,N+3Dataset i=0,,N+4i=0,...,N+4Initial set i=0,,Ni=0,...,NBatch Rolling WindowDataset i=1,,N+1i=1,...,N+1Dataset i=2,,N+2i=2,...,N+2Dataset i=3,,N+3i=3,...,N+3Dataset i=4,,N+4i=4,...,N+4Initial set i=0,,Ni=0,...,NOnline LearningNN+1NN+2New observations added row-by-row.NN+3NN+4
@@
We use the same data set as in Lipiecki et al. (2024), which consists of electricity prices for the German day-ahead market from 2015-01-01 to 2024-01-01 exclusive. In line with previous works, we use the data until 2018-12-26 as initial training set, leaving 1831 observations and therefore more than 4 years, as it is best practice (Lago et al., 2021), for out-of-sample testing. This dataset has been employed in various studies and to enable comparability, we further split the dataset in two sub-samples, that is 2018-12-27 to 2020-12-31 (736 days) and 2021-01-01 to 2023-12-31 (1095 days). Additionally, the data set contains day-ahead renewable production forecasts, load forecasts and prices for fundamental commodities. All features are briefly described in Table 3. We use an incremental mean-variance scaling to post-process the electricity prices. We denote the electricity price for day $t$ and hour $h\in 0,...,23$ as $P_{th}$. We have $y_{th}=(P_{th}-\tilde{\mu}_{th})/\tilde{\sigma}_{th}$, where $\tilde{\mu}_{th}=1/t\sum_{i}^{t}P_{ih}$ and $\tilde{\sigma}_{th}=\sqrt{1/t\sum_{i}^{t}(P_{ih}-\tilde{\mu}_{ih})^{2}}$ are the mean and standard deviation up to observation $t,h$ and therefore have $\bm{\mathrm{Y}}=(\bm{\mathrm{y}}_{0},...,\bm{\mathrm{y}}_{23})$ as the 24-dimensional ($D=H=24$) response matrix. We find that this stabilizes the estimation of covariance matrices and the normalization can be re-applied after the estimation, i.e. the covariance of $\bm{\mathrm{p}}_{t}$ corresponds to $\operatorname{diag}(\tilde{\bm{\mathrm{\sigma}}}_{t})\bm{\mathrm{\Sigma}}\operatorname{diag}(\tilde{\bm{\mathrm{\sigma}}}_{t})$, where $\bm{\mathrm{\Sigma}}$ is estimated based on $\bm{\mathrm{y}}_{t}$. Incremental updates of the mean and variance are straight-forward in an online learning setting using Welford’s method (Welford, 1962). Note that, in this application study, we have $T$ corresponding to $N$ and $H$ corresponding to $D$ in the general notation.
@@
$\operatorname{ResLoad}_{t,h}$ Day-ahead residual load forecast Hourly ENTSO-E
$\overline{\operatorname{ResLoad}}_{t}$ Day-ahead baseload residual load forecast $\frac{1}{H}\sum_{h=1}^{H}\operatorname{ResLoad}_{t,h}$ Daily ENTSO-E
@@
Table 3: Variables from the data set of Marcjasz et al. (2023) and Lipiecki et al. (2024).
3.2 Model Definition and Benchmarks